thrupti:

Calculate the pH of the resulting solution if 34.0 mL of 0.340 M HCl(aq) is added to 44.0 mL of 0.340 M NaOH(aq).

Vocaloid:

the molarities are the same which makes things a bit simpler the 34.0 mL of HCl is neutralized by 34.0 mL of NaOH, leaving 44.0 - 34.0 mL = 10.0 mL of NaOH leftover inside 34.0 + 34.0 = 68.0 mL of water, resulting in 10.0 + 68.0 = 78.0 mL of solution calculate the moles of NaOH for 10.0 mL at a molarity of 0.340 M, divide by the total L of solution, then -log[NaOH] gives pOH which you can then simply find 14 - pOH to find pH